### v5

@@ Abstract
Defining pathologies automatically from medical images aids the understanding of the emergence and progression of diseases, and such an ability is crucial in clinical diagnostics. However, existing deep learning models heavily rely on expert annotations and lack generalization capabilities in open clinical environments. In this study, we present a generalizable vision-language model for Annotation-Free pathology Localization (AFLoc). The core strength of AFLoc lies in its extensive multi-level semantic structure-based contrastive learning, which comprehensively aligns multi-granularity medical concepts from reports with abundant image features, to adapt to the diverse expressions of pathologies and unseen pathologies without the reliance on image annotations from experts. We conducted primary experiments on a dataset of 220K pairs of image-report chest X-ray images, and performed extensive validation across six external datasets encompassing 20 types of chest pathologies. The results demonstrate that AFLoc outperforms state-of-the-art methods in both annotation-free localization and classification tasks. Additionally, we assessed the generalizability of AFLoc on other modalities, including histopathology and retinal fundus images. Extensive experiments show that AFLoc exhibits robust generalization capabilities, even surpassing human benchmarks in localizing five different types of pathological images. These results highlight the potential of AFLoc in reducing annotation requirements and its applicability in complex clinical environments.
@@
Accurate diagnosis and precise pathology localization in medical images are paramount for optimizing treatment effectiveness and interventions. These technologies are instrumental in personalized medicine, facilitating customized treatment approaches that improve patient outcomes and mitigate the possibility of diagnostic errors. By pinpointing the exact location and extent of abnormalities, clinicians can make informed decisions that lead to more targeted therapies and improved prognoses for patients [1, 2, 3, 4].
@@
Over the past decade, supervised deep learning methods have significantly accelerated advancements in disease localization [5, 6, 7, 8]. However, the efficacy of these methods heavily relies on extensively annotated training datasets, which require domain experts to invest considerable time [9]. Specifically, clinical localization tasks often require experienced clinicians to meticulously annotate numerous precise bounding boxes or perform pixel-wise delineations of localized pathology areas. This annotation process is costly, particularly in resource-constrained clinical settings, and algorithms frequently struggle to generalize to diverse datasets.
@@
Several methods have been proposed to reduce the reliance on large annotated datasets [10, 11, 12, 13]. Initially, these methodologies acquire general visual representations through self-supervised learning from image datasets, followed by fine-tuning on smaller annotated datasets. This approach enables models to achieve high performance on specific tasks while significantly decreasing the need and cost of data labeling [13]. Moreover, saliency-based methods [14, 15, 16, 17] have been developed to reduce annotation costs in pathology localization tasks by allowing coarse localization of target categories in models trained with image-level annotations. However, these methods still require annotations for specific downstream tasks. This requirement is particularly challenging in flexible and dynamic clinical environments, especially for newly emerging diseases (e.g., COVID-19), where deployed models may fail to perform effectively [18, 19, 20].
@@
In recent years, unsupervised deep learning methods have gained increasing attention due to their independence from annotated datasets, particularly in the field of anomaly detection [21, 22, 23, 24]. These methods typically train models using only healthy samples, learning the distribution of normal anatomical structures, which enables the identification of abnormal pathology samples during the testing phase [21, 22]. They are particularly effective for data with simple structures and low inter-sample variance, allowing them to learn normative distributions and achieve excellent anomaly detection performance [23, 24, 25]. However, challenges such as the high heterogeneity of pathology images, significant similarities between different pathologies, and large variations in contrast for the same lesions reduce the usability of these methods in complex scenarios, thereby hindering their practical application in real medical environments [25].
@@
A promising approach is the development of medical visual-language pre-training methods [26, 27, 28, 29, 19, 30, 31]. These methods establish effective correlations between medical reports and medical images, allowing them to flexibly localize disease types not encountered during pre-training without requiring additional customized annotations [19]. However, achieving precise pathology localization solely through the combination of medical images and clinical reports remains challenging. A primary obstacle is the lack of explicit pathology localization markers in clinical reports, which often provide only coarse information such as ‘upper’ or ‘left’ to indicate disease location. Moreover, clinical descriptions by clinicians are subjective and variable, further complicating the task of accurately extracting and localizing diseases in medical images. To address this challenge, several methods have been proposed to integrate finer-grained information. For instance, GLoRIA [30] extracts the correlation of the image’s regions and paired words in reports to learn global and local representations of images. MedKLIP [19] utilizes well-defined medical vocabulary knowledge bases to provide supervision at the entity level through triplet training paradigms. However, these fine-grained methods typically focus on individual levels of medical concepts and may overlook the variable meanings of concepts in different contexts. Therefore, these approaches may struggle to adapt to the diverse expressions of disease descriptors in clinical practice, often requiring customized textual cues to enhance localization performance.
@@
In this study, we propose AFLoc, a visual-language model based on contrastive learning aimed at alleviating the need for costly pathology localization annotations. AFLoc can autonomously perform pathology localization and clinical diagnosis with medical images. Unlike traditional global semantic alignment strategies [29, 28], AFLoc introduces a contrastive learning framework with a multi-level semantic alignment component, facilitating the comprehensive alignment of medical concepts from reports with image features. Specifically, the image encoder generates three levels of features: shallow local features, deep local features, and global features, which are aligned with word-level, sentence-level, and report-level features extracted by the text encoder. We extensively validated AFLoc across three types of medical image datasets, including Chest X-Ray (6 external datasets), histopathology (3 external datasets), and retinal fundus images. Our results show that AFLoc outperforms state-of-the-art methods in localization and clinical diagnostic tasks across different modalities. We hope that this study can help address the challenges posed by annotation scarcity and modality diversity in clinical environments, while providing new insights for the design of future clinical open-environment methods.
a
b
@@ Results
AFLoc is capable of performing zero-shot localization tasks without requiring any annotations. In this section, we present AFLoc’s annotation-free localization performance across different clinical modalities. Each modality has been pre-trained using its corresponding image-report datasets, ensuring modality-specific knowledge is effectively captured and utilized. Here, we consider that ’pathology localization’ encompasses both heatmap predictions and binary mask results, as some clinical applications may benefit more from the visualization provided by heatmaps rather than discrete segmentation. We quantitatively evaluate each predicted heatmap using Contrast-to-Noise Ratio (CNR) [28] and assess the thresholded segmentation masks using Intersection over Union (IoU) [32] and Dice similarity coefficient metrics. In the following, we will present the results of the three modalities separately.
@@
Localization Performance on Chest X-Ray In clinical practice for chest diseases, doctors often need to perform detailed assessments of Chest X-rays for early screening and surgical planning. For instance, airspace opacity can indicate lung infections, inflammation, or tumors [33]. Atelectasis leads to restricted gas exchange, causing hypoxemia and severely affecting respiratory function [34]. However, since these diseases often lack clear boundaries and have contrast levels similar to surrounding tissues, doctors typically need to invest significant effort in distinguishing them. Therefore, automated diagnostic systems play a crucial role in aiding diagnosis.
@@
We evaluated the localization performance of AFLoc on Chest X-rays using four external datasets: RSNA Pneumonia [35], MS-CXR [28], CheXlocalize [14], and COVID Rural [36, 37]. These datasets cover 13 common thoracic pathologies: pneumonia, airspace opacity, atelectasis, cardiomegaly, consolidation, edema, enlarged cardiomediastinum, lung lesion, pleural effusion, pneumothorax, lung opacity, support devices, and COVID-19.
@@
We first elaborate AFLoc’s localization performance on the RSNA Pneumonia dataset (Fig. 2, Extended Data Table 3). AFLoc achieves better results than all comparative methods across all three evaluation metrics. Specifically, compared to the respective best-performing comparative methods, AFLoc boosts the IoU by 7.9% from 0.317 (95% CI: 0.309, 0.325) to 0.342 (95% CI: 0.332, 0.350), enhances the Dice coefficient by 4.1% from 0.465 (95% CI: 0.455, 0.474) to 0.484 (95% CI: 0.474, 0.495), and elevates CNR by 18.3% from 1.300 (95% CI: 1.262, 1.339) to 1.538 (95% CI: 1.496, 1.580). These results highlight AFLoc’s superior ability to localize pathological regions.
@@
Similarly, the results on the MS-CXR dataset (Fig. 2, Extended Data Table 6) show that our AFLoc outperforms the existing vision-language pre-training approaches (GLoRIA, BioViL, and MedKLIP) by a clear margin. For example, AFLoc improves the IoU (0.324, 95% CI: 0.298, 0.350) by more than 6% compared to GLoRIA (0.268, 95% CI: 0.242, 0.295), BioViL (0.228, 95% CI: 0.201, 0.255), and MedKLIP (0.264, 95% CI: 0.240, 0.288). In addition, AFLoc shows consistent performance enhancement across the majority of pathologies (six out of eight) (Fig. 3), confirming its superior pathology localization capability. We also compared AFLoc with saliency-based methods (Grad-CAM, Grad-CAM++, Eigen-CAM, and the human benchmark) on this dataset using published results from CheXlocalize (Fig. 3b). AFLoc outperforms these methods in most cases, achieving a 4% higher average IoU than Grad-CAM (0.318 vs. 0.282). AFLoc also surpasses the human benchmark in localizing five pathologies, including airspace opacity and pleural effusion, demonstrating its potential in medical image analysis.
@@
Localization Performance on Retinal Fundus Images To further demonstrate the generalizability of our AFLoc, we applied it to retinal fundus images. We curated a retinal fundus image dataset containing three prevalent retinal pathologies: choroidal neovascularization (CNV), drusen, and intraretinal hemorrhages. Detecting choroidal neovascularization is essential for the early diagnosis of wet age-related macular degeneration (AMD) [38]. Drusen serves as an indicator of dry AMD [38]. Intraretinal hemorrhages can reflect severe conditions such as diabetic retinopathy or hypertensive retinopathy [39]. Early detection and identification of these pathologies can significantly slow disease progression and facilitate proactive interventions.
The localization results of different methods on these three retinal pathologies are depicted in Fig. 2 (column 4) and Extended Data Table 4. Notably, AFLoc attains the highest IoU, Dice, and CNR scores across the different pathologies. For choroidal neovascularization, AFLoc achieves an IoU of 0.321 (95% CI: 0.303, 0.337), which is much higher than GLoRIA (0.143, 95% CI: 0.130, 0.156) and other comparative methods. Regarding drusen, AFLoc achieves an IoU of 0.147 (95% CI: 0.125, 0.170) and a Dice score of 0.225 (95% CI: 0.192, 0.257), showcasing its proficiency in localizing subtle features. While AFLoc achieves superior scores compared to all comparative methods, its capability to localize intraretinal hemorrhages is relatively weaker than for the other two pathologies, with an IoU of 0.097 (95% CI: 0.091, 0.102). This suggests the necessity for further enhancements in localizing diffuse and less distinct lesions. In Fig. 4, the visualization of retinal localization is presented, with deeper red areas denoting model-predicted disease locations, while the black box signifies clinician annotations. AFLoc accurately localizes concentrated pathologies like choroidal neovascularization and drusen.
@@
[…]
Here, we utilized the Quilt-1M dataset [41] to train the AFLoc model for histopathology and tested its performance on the SICAPv2 test set (n=2,122) under the annotation-free localization setting. We compared AFLoc with several state-of-the-art methods, including CLIP [42], BiomedCLIP [43], and three foundation models specialized for histopathology: PLIP [44], QuiltNet [41], and CONCH [45]. Specifically, we applied GradCAM [15] for the comparative methods, as it has been shown to enhance their performance.
As shown in Fig. 2 and Extended Data Table 5, AFLoc achieves the best localization performance across all three evaluation metrics. Specifically, compared to the best-performing comparative method, PLIP, AFLoc demonstrates notable improvements in IoU (0.282, 95% CI: 0.264, 0.300 vs. 0.249, 95% CI: 0.230, 0.269), Dice score (0.396, 95% CI: 0.375, 0.418 vs. 0.352, 95% CI: 0.327, 0.376), and CNR (0.639, 95% CI: 0.572, 0.706 vs. 0.321, 95% CI: 0.249, 0.398). These results suggest that AFLoc is capable of more accurately localizing abnormal tissues compared to both general-purpose and histopathology-specific models.
@@
To explore this, we compared the results obtained using simple and precise prompts on the MS-CXR dataset. For example, the simple prompt for pneumonia can be “findings suggesting pneumonia”, while the precise description can be “severe bibasilar consolidation” or “air space opacity in a right infrahilar location”, reflecting clinical terminology.
Results in Table 1 show that when utilizing precise prompts, all models, including our AFLoc and the two comparative methods (BioViL and GLoRIA), demonstrate enhanced localization performance. For our AFLoc, the utilization of precise prompts leads to a 12.1% increase in IoU (from 0.289 to 0.324), a 10.5% increase in Dice (from 0.418 to 0.462), and a 21.1% increase in CNR (from 1.351 to 1.636). These improvements highlight the advantages of detailed, clinically relevant prompts to enhance localization accuracy. Nevertheless, AFLoc outperforms the two comparative methods regardless of the prompt generation methods.
Fig. 5 illustrates that with simple prompts, models tend to have broader, less accurate focal areas, frequently failing to align with radiologist-annotated regions. In contrast, with precise prompts, AFLoc exhibits enhanced focus, closely matching the annotations made by radiologists. For example, in the case of left lower lobe pneumonia, AFLoc accurately highlights the affected region with precise prompts. Similarly, for the case involving tiny right basilar pneumothorax, much better localization performance is observed when precise prompts are utilized compared to simple prompts.
| Description | BioViL | GLoRIA | AFLoc |
| IoU | |||
| Simple description | 0.187 (0.162,0.213) | 0.240 (0.214,0.265) | 0.289 (0.262,0.314) |
| Precise description | 0.228 (0.201,0.255) | 0.268 (0.242,0.295) | 0.324 (0.298,0.350) |
| Dice | |||
| Simple description | 0.287 (0.252,0.321) | 0.357 (0.322,0.392) | 0.418 (0.383,0.450) |
| Precise description | 0.342 (0.308,0.377) | 0.392 (0.357,0.427) | 0.462 (0.431,0.492) |
| CNR | |||
| Simple description | 0.825 (0.694,0.948) | 1.128 (1.003,1.254) | 1.351 (1.219,1.481) |
| Precise description | 1.083 (0.949,1.219) | 1.287 (1.149,1.421) | 1.636 (1.501,1.772) |
@@
In this section, we demonstrate AFLoc’s annotation-free diagnostic performance through zero-shot classification, using CXR, retinal fundus images, and histopathology datasets. For comparison, we also present the results of two unsupervised anomaly detection methods, ReContrast and RD++. These methods, without fine-tuning using labeled datasets, are referred to as “zero-shot” methods for convenience.
[…]
Diagnosis Performance on Retinal Fundus Images We further applied AFLoc to retinal fundus images, the results are presented in Table 2. The evaluation covered nine retinal diseases: Macular Degeneration (MD), Retinopathy, Myopia, Glaucoma, Congenital Optic Disc Anomalies (CODA), Retinal Arteriosclerosis (RAS), Cataract, Macular Epiretinal Membrane (MEM), and Macular Lesion (ML). While unsupervised anomaly detection methods, ReContrast and RD++, perform well on specific diseases like Myopia (0.921 AUROC) and Cataract (0.819 AUROC), their average scores are lower (0.589 and 0.52, respectively) compared to vision-language pre-training methods. Among the label free methods, AFLoc achieves the highest scores across all diseases, with an average AUROC of 0.908, much higher than the second-best method, GLoRIA (0.772). Additionally, AFLoc’s performance is comparable to RETFound, which used labeled data for fine-tuning. These results highlight AFLoc’s exceptional diagnostic capability and its potential for generalizing across diverse diseases.
[…]
a
b

### v7

@@ Abstract
Existing deep learning models for defining pathology from clinical imaging data rely on expert annotations and lack generalization capabilities in open clinical environments. Here, we present a generalizable vision-language model for Annotation-Free pathology Localization (AFLoc). The core strength of AFLoc is extensive multi-level semantic structure-based contrastive learning, which comprehensively aligns multi-granularity medical concepts with abundant image features to adapt to the diverse expressions of pathologies without the reliance on expert image annotations. We conduct primary experiments on a dataset of 220K pairs of image-report chest X-ray images and perform validation across eight external datasets encompassing 34 types of chest pathologies. The results demonstrate that AFLoc outperforms state-of-the-art methods in both annotation-free localization and classification tasks. Additionally, we assess the generalizability of AFLoc on other modalities, including histopathology and retinal fundus images. We show that AFLoc exhibits robust generalization capabilities, even surpassing human benchmarks in localizing five different types of pathological images. These results highlight the potential of AFLoc in reducing annotation requirements and its applicability in complex clinical environments.
@@
Accurate diagnosis and precise pathology localization in medical images facilitate customized treatment approaches that improve patient outcomes and mitigate the possibility of diagnostic errors. By pinpointing the exact location and extent of abnormalities, clinicians can make informed decisions that lead to more targeted therapies and improved prognoses for patients [coudray2018classification, wang2018development, campanella2019clinical, courtiol2019deep].
@@
Over the past decade, supervised deep learning methods have accelerated advancements in disease localization [lee2022localization, cao2023large, song2020clinically, leon2023hyperspectral]. However, the efficacy of these methods heavily relies on extensively annotated training datasets, which require domain experts to invest considerable time [rajpurkar2022ai, wang_annotation-efficient_2021]. Specifically, clinical localization tasks often require experienced clinicians to meticulously annotate numerous precise bounding boxes or perform pixel-wise delineations of localized pathology areas. This annotation process is costly, particularly in resource-constrained clinical settings, and algorithms frequently struggle to generalize to diverse datasets.
@@
Several methods have been proposed to reduce the reliance on large annotated datasets [wang_annotation-efficient_2021, liu2024swin, chen2020simple, zhou2023foundation]. Initially, these methodologies acquire general visual representations through self-supervised learning from image datasets, followed by fine-tuning on smaller annotated datasets. This approach enables models to achieve high performance on specific tasks while decreasing the need and cost of data labeling [zhou2023foundation]. Moreover, saliency-based methods [saporta2022benchmarking, gradcam, gradcam++, eigencam] have been developed to reduce annotation costs in pathology localization tasks by allowing coarse localization of target categories in models trained with image-level annotations. However, these methods still require annotations for specific downstream tasks. This requirement is particularly challenging in flexible and dynamic clinical environments, especially for emerging diseases (e.g., COVID-19), where deployed models may fail to perform effectively [tiu2022expert, wu2023medklip, morens2013emerging].
@@
In recent years, unsupervised deep learning methods have gained increasing attention due to their independence from annotated datasets, particularly in the field of anomaly detection [baur2021autoencoders, rd, rd++, ReContrast]. These methods typically train models using only healthy samples, learning the distribution of normal anatomical structures, which enables the identification of abnormal pathology samples during the testing phase [baur2021autoencoders, rd]. They are particularly effective for data with simple structures and low inter-sample variance, allowing them to learn normative distributions and achieve excellent anomaly detection performance [rd++, ReContrast, upd]. However, challenges such as the high heterogeneity of pathology images, similarities between different pathologies, and large variations in contrast for the same lesions reduce the usability of these methods in complex scenarios, thereby hindering their practical application in real medical environments [upd].
@@
A promising approach is the development of medical visual-language pre-training methods [huang_nc_2024, huang_mia_2024, boecking2022making, convirt, wu2023medklip, huang2021gloria, zhou2023advancing]. These methods establish effective correlations between medical reports and medical images, allowing them to flexibly localize disease types not encountered during pre-training without requiring additional customized annotations [wu2023medklip]. However, achieving precise pathology localization solely through the combination of medical images and clinical reports remains challenging. A primary obstacle is the lack of explicit pathology localization markers in clinical reports, which often provide only coarse information such as ‘upper’ or ‘left’ to indicate disease location. Moreover, clinical descriptions by clinicians are subjective and variable, further complicating the task of accurately extracting and localizing diseases in medical images. To address this challenge, several methods have been proposed to integrate finer-grained information. For instance, GLoRIA [huang2021gloria] extracts the correlation of the image’s regions and paired words in reports to learn global and local representations of images. MedKLIP [wu2023medklip] uses well-defined medical vocabulary knowledge bases to provide supervision at the entity level through triplet training paradigms. However, these fine-grained methods typically focus on individual levels of medical concepts and may overlook the variable meanings of concepts in different contexts. Therefore, these approaches may struggle to adapt to the diverse expressions of disease descriptors in clinical practice, often requiring customized textual cues to enhance localization performance.
@@
In this study, we propose AFLoc, a visual-language model based on contrastive learning aimed at alleviating the need for costly pathology localization annotations. AFLoc can autonomously perform pathology localization and clinical diagnosis with medical images. Unlike traditional global semantic alignment strategies [convirt, boecking2022making], AFLoc introduces a contrastive learning framework with a multi-level semantic alignment component, facilitating the comprehensive alignment of medical concepts from reports with image features. Specifically, the image encoder generates three levels of features: shallow local features, deep local features, and global features, which are aligned with word-level, sentence-level, and report-level features extracted by the text encoder. We extensively validated AFLoc across three types of medical image datasets, including chest X-ray (8 external datasets), histopathology (3 external datasets), and retinal fundus images. Our results show that AFLoc outperforms state-of-the-art methods in localization and clinical diagnostic tasks across different modalities. We hope that this study can help address the challenges posed by annotation scarcity and modality diversity in clinical environments, while providing insights for the design of future clinical open-environment methods.
@@ Results
AFLoc is capable of performing zero-shot localization tasks without requiring any annotations. In this section, we present AFLoc’s annotation-free localization performance across different clinical modalities. Each modality has been pre-trained using its corresponding image-report datasets, ensuring modality-specific knowledge is effectively captured and used. Here, we consider that ’pathology localization’ encompasses both heatmap predictions and binary mask results, as some clinical applications may benefit more from the visualization provided by heatmaps rather than discrete segmentation. We quantitatively evaluate each predicted heatmap using Contrast-to-Noise Ratio (CNR) [boecking2022making] and assess the thresholded segmentation masks using Intersection over Union (IoU) [rezatofighi2019generalized] and Dice similarity coefficient metrics. In the following, we will present the results of the three modalities separately.
@@
Localization Performance on Chest X-Ray In clinical practice for chest diseases, doctors often need to perform detailed assessments of chest X-rays for early screening and surgical planning. For instance, airspace opacity can indicate lung infections, inflammation, or tumors [reed2011multifocal]. Atelectasis leads to restricted gas exchange, causing hypoxemia and severely affecting respiratory function [woodring1996types]. However, since these diseases often lack clear boundaries and have contrast levels similar to surrounding tissues, doctors typically need to invest significant effort in distinguishing them. Therefore, automated diagnostic systems play a crucial role in aiding diagnosis.
@@
We evaluated the localization performance of AFLoc on chest X-rays using four external datasets: RSNA Pneumonia [shih2019augmenting], MS-CXR [boecking2022making], CheXlocalize [saporta2022benchmarking], and COVID Rural [desai2020chest, tang2020deep]. These datasets cover 13 common thoracic pathologies: pneumonia, airspace opacity, atelectasis, cardiomegaly, consolidation, edema, enlarged cardiomediastinum, lung lesion, pleural effusion, pneumothorax, lung opacity, support devices, and COVID-19.
@@
We first elaborate AFLoc’s localization performance on the RSNA Pneumonia dataset (Fig. 2, Supplementary Table3). AFLoc achieves better results than all comparative methods across all three evaluation metrics. Specifically, compared to the respective best-performing comparative methods, AFLoc boosts the IoU by 7.9% from 0.317 (95% CI: 0.309, 0.325) to 0.342 (95% CI: 0.332, 0.350), enhances the Dice coefficient by 4.1% from 0.465 (95% CI: 0.455, 0.474) to 0.484 (95% CI: 0.474, 0.495), and elevates CNR by 18.3% from 1.300 (95% CI: 1.262, 1.339) to 1.538 (95% CI: 1.496, 1.580). These results highlight AFLoc’s superior ability to localize pathological regions.
@@
Similarly, the results on the MS-CXR dataset (Fig. 2, Supplementary Table3 and 6) show that our AFLoc outperforms the existing vision-language pre-training approaches (GLoRIA, BioViL, and MedKLIP) by a clear margin. For example, AFLoc improves the IoU (0.324, 95% CI: 0.298, 0.350) by more than 6% compared to GLoRIA (0.268, 95% CI: 0.242, 0.295), BioViL (0.228, 95% CI: 0.201, 0.255), and MedKLIP (0.264, 95% CI: 0.240, 0.288). AFLoc shows consistent performance enhancement across the majority of pathologies (six out of eight) (Fig. 3a), confirming its superior pathology localization capability. In addition, on the CheXlocalize dataset, we compared AFLoc with saliency-based methods (Grad-CAM, Grad-CAM++, and Eigen-CAM) and the human benchmark using results reported in the literature [saporta2022benchmarking] (Fig. 3b). AFLoc outperforms these methods in most cases, achieving a 4% higher average IoU than Grad-CAM (0.318 vs. 0.282). AFLoc also surpasses the human benchmark in localizing five pathologies, including airspace opacity and pleural effusion, demonstrating its potential in medical image analysis.
@@
Localization Performance on Retinal Fundus Images To further demonstrate the generalizability of our AFLoc, we applied it to retinal fundus images. We curated a retinal fundus image dataset containing three prevalent retinal pathologies: choroidal neovascularization (CNV), drusen, and intraretinal hemorrhages. Detecting choroidal neovascularization is essential for the early diagnosis of wet age-related macular degeneration (AMD) [lim2012age]. Drusen serves as an indicator of dry AMD [lim2012age]. Intraretinal hemorrhages can reflect severe conditions such as diabetic retinopathy or hypertensive retinopathy [uhler2008optic]. Early detection and identification of these pathologies can significantly slow disease progression and facilitate proactive interventions.
The localization results of different methods on these three retinal pathologies are depicted in Fig. 2 (column 4) and Supplementary Table4. Notably, AFLoc attains the highest IoU, Dice, and CNR scores across the different pathologies. For choroidal neovascularization, AFLoc achieves an IoU of 0.321 (95% CI: 0.303, 0.337), which is much higher than GLoRIA (0.143, 95% CI: 0.130, 0.156) and other comparative methods. Regarding drusen, AFLoc achieves an IoU of 0.147 (95% CI: 0.125, 0.170) and a Dice score of 0.225 (95% CI: 0.192, 0.257), showcasing its proficiency in localizing subtle features.
While AFLoc achieves superior scores compared to all comparative methods, its capability to localize intraretinal hemorrhages is relatively weaker than for the other two pathologies, with an IoU of 0.097 (95% CI: 0.091, 0.102). This suggests the necessity for further enhancements in localizing diffuse and less distinct lesions. In Fig. 4, the visualization of retinal localization is presented, with deeper red areas denoting model-predicted disease locations, while the black box signifies clinician annotations. AFLoc accurately localizes concentrated pathologies like choroidal neovascularization and drusen.
@@
Localization Performance on Histopathology In histopathology, accurate localization of abnormal tissues is key for cancer diagnosis and grading. For example, in prostate cancer, the Gleason grading system is used to assess tissue samples by examining glandular architecture, nuclear features, and lumen formation. Pathologists assign a Gleason score based on the tissue’s differentiation, reflecting the cancer’s aggressiveness [sicap]. Early detection and identification of these pathologies can significantly slow disease progression and facilitate preventive interventions.
Here, we used the Quilt-1M dataset [quilt] to train the AFLoc model for histopathology and tested its performance on the SICAPv2 test set (n=2,122) under the annotation-free localization setting. We compared AFLoc with several state-of-the-art methods, including CLIP [clip], BiomedCLIP [biomedclip], and three foundation models specialized for histopathology: PLIP [plip], QuiltNet [quilt], and CONCH [conch]. Specifically, we applied GradCAM [gradcam] for the comparative methods, as it has been shown to enhance their performance.
As shown in Fig. 2 and Supplementary Table5, AFLoc achieves the best localization performance across all three evaluation metrics. Specifically, compared to the best-performing comparative method, PLIP, AFLoc demonstrates notable improvements in IoU (0.285, 95% CI: 0.267, 0.304 vs. 0.249, 95% CI: 0.230, 0.269), Dice score (0.399, 95% CI: 0.378, 0.421 vs. 0.352, 95% CI: 0.327, 0.376), and CNR (0.630, 95% CI: 0.561, 0.697 vs. 0.321, 95% CI: 0.249, 0.398). These results suggest that AFLoc is capable of more accurately localizing abnormal tissues compared to both general-purpose and histopathology-specific models.
@@
To explore this, we compared the results obtained using simple and precise prompts on the MS-CXR dataset. For example, the simple prompt for pneumonia can be ‘findings suggesting pneumonia’, while the precise description can be ‘severe bibasilar consolidation’ or ‘air space opacity in a right infrahilar location’, reflecting clinical terminology.
Results in Table 1 show that when using precise prompts, all models, including our AFLoc and the two comparative methods (BioViL and GLoRIA), demonstrate enhanced localization performance. For our AFLoc, the use of precise prompts leads to a 12.1% increase in IoU (from 0.289 to 0.324), a 10.5% increase in Dice (from 0.418 to 0.462), and a 21.1% increase in CNR (from 1.351 to 1.636). These improvements highlight the advantages of detailed, clinically relevant prompts to enhance localization accuracy. Nevertheless, AFLoc outperforms the two comparative methods regardless of the prompt generation methods.
Fig. 5 illustrates that with simple prompts, models tend to have broader, less accurate focal areas, frequently failing to align with radiologist-annotated regions. In contrast, with precise prompts, AFLoc exhibits enhanced focus, closely matching the annotations made by radiologists. For example, in the case of left lower lobe pneumonia, AFLoc accurately highlights the affected region with precise prompts. Similarly, for the case involving tiny right basilar pneumothorax, much better localization performance is observed when precise prompts are used compared to simple prompts. Moreover, we show the model’s ability to generate distinct localization responses when prompted with different disease descriptions, as shown in Extended Data Fig. 8. These examples include multi-lesion cases with varying degrees of spatial overlap. Notably, the model accurately localizes the corresponding abnormality based on the prompts with simple descriptions for different diseases.
| Description | BioViL | GLoRIA | AFLoc | ||||||
| IoU | |||||||||
| Simple description |
| Precise description |
| Dice | |||||||||
| Simple description |
| Precise description |
| CNR | |||||||||
| Simple description |
| Precise description |
Bold values represent the highest performance score among the compared methods.
@@
In this section, we demonstrate AFLoc’s annotation-free diagnostic performance through zero-shot classification, using CXR, retinal fundus images, and histopathology datasets. For comparison, we also present the results of two unsupervised anomaly detection methods, ReContrast and RD++. These methods, without fine-tuning using labeled datasets, are referred to as ‘zero-shot’ methods for convenience.
Diagnosis Performance on Chest X-Ray In our study, we conducted zero-shot classification tasks on the RSNA Pneumonia, SIMM, NIH ChestXray14, and CXR-LT datasets. As depicted in Fig. 6a, the saliency-based methods (ReContrast and RD++) yield lower scores in comparison to the vision-language multi-modal pre-training techniques. This highlights the significance of incorporating expert-knowledge-rich text reports. MedKLIP excels in this task, likely due to its distinctive semantic simplification of text reports, facilitating effective learning of classification-relevant information. Nevertheless, our AFLoc achieves the highest AUROC scores of 0.881 (RSNA), 0.902 (SIIM), 0.737 (NIH ChestXray14), 0.735 (CXR-LT Task 2), and 0.726 (CXR-LT Task 3) on the four datasets, respectively. Supplementary Table7 provides detailed comparisons across various datasets and diseases. AFLoc demonstrates notable performance improvements over MedKLIP for specific diseases like atelectasis, pneumothorax, and pleural thickening. These results underscore the AFLoc’s ability to effectively align multi-modal information, leveraging both the visual and textual features to capture complex classification patterns.
Diagnosis Performance on Retinal Fundus Images We further applied AFLoc to retinal fundus images, the results are presented in Table 2. The evaluation covered nine retinal diseases: Macular Degeneration (MD), Retinopathy, Myopia, Glaucoma, Congenital Optic Disc Anomalies (CODA), Retinal Arteriosclerosis (RAS), Cataract, Macular Epiretinal Membrane (MEM), and Macular Lesion (ML). While unsupervised anomaly detection methods, ReContrast and RD++, perform well on specific diseases like Myopia (0.921 AUROC) and Cataract (0.819 AUROC), their average scores are lower (0.589 and 0.520, respectively) compared to vision-language pre-training methods. Among the label free methods, AFLoc achieves the highest scores across all diseases with an average AUROC of 0.908, which is much higher than the second-best method GLoRIA (0.772). Additionally, AFLoc’s performance is comparable to RETFound, which used labeled data for fine-tuning. These results highlight AFLoc’s exceptional diagnostic capability and its potential for generalizing across diverse diseases.
Diagnosis Performance on Histopathology The gold standard for diagnosing many diseases remains the examination of histopathology images. Supplementary Table8 and Fig. 6b present the comparison of balanced accuracy for various methods on the SICAPv2, WSSS4LUAD, and DHMC LUAD datasets for the zero-shot classification task. Among the compared methods, AFLoc demonstrates superior performance across all datasets, achieving the highest balanced accuracy scores of 0.512, 0.704, and 0.442 on the SICAPv2, WSSS4LUAD, and DHMC LUAD datasets, respectively. The consistent performance enhancement of AFLoc can be attributed to its robust multi-modal semantic alignment, which effectively captures both visual features from histopathology images and semantic context from textual descriptions.
@@
Bold values represent the highest performance score among the compared methods.